# Learning the symmetric group: large from small

### v1

@@ 1.1 Related work on symmetric group
The concept of grokking was introduced in Power2022 where the authors considered simple algorithmic datasets and showed that validation accuracy increases towards perfect generalization long after the training accuracy becomes close to perfect. Despite overfitting early in the training process, continued optimization eventually leads to a transition where the model generalizes. This suggests that the network gradually shifts from memorization to learning the underlying algorithm. One example from Power2022 is grokking of group multiplication in the permutation group $S_{5}$. For a given pair of permutations $a$ and $b$ the model needs to compute a permutation $c$ such that $a\circ b=c$. Grokking and interpretability of the group multiplication in small permutation groups has been studied in a series of subsequent works ChughtaiCN2021 ; StanderYFB2023 ; Wu2024 (see also NandaCLSS2023 ). In the present work we also consider a symmetric group problem but in contrast with previous works we ask whether a model is able to generalize from training on $S_{k}$ data to testing on $S_{n}$ data with $k\ll n$. Although we have not observed grokking, test accuracy is nearly 100%. At the time of writing it is an open question whether our models have discovered a general algorithm.
Our result is reminiscent of out of distribution (OOD) learning. In mathematical problems OOD learning has been studied in, for example, Charton2022 . We test OOD learning in the context of the symmetric group, using a scalable presentation with general transpositions as group generators, as well as a less scalable, more local approach using only elementary transpositions. One motivation for this study is to identify an effective machine learning approach to improve on the unknot problem GukovHRS:2020 .
An important part of earlier works on grokking in symmetric group is analysis of models using mechanistic interpretability. This can be done using different circuit level approaches Olah2020 ; Elhage2021 ; Olsson2022 ; Quirke2024 ; Zhang2023 ; Zhong2023 ; Gross2024 . Examining activations of individual neurons and intervening in various parts of the network can help us to discover circuits that implement the algorithm. We aim to perform a similar analysis of our models in a future work.
@@ 3 What are we learning?
The experiment reported in this note investigates whether a transformer can learn the symmetric group $S_{n}$ by training only on subgroups isomorphic to $S_{m}$ with $m<n$. That is, given a word $w\in S_{n}$ can the corresponding permutation $(\sigma(1),\ldots,\sigma(n)):=(1,\ldots,n)\cdot w$ be correctly predicted by a transformer without explicitly hardcoding the group relations (4).
@@ 3.1 General transpositions
$\Phi:\mathbf{x}\mapsto\mathbf{p},$
@@
In order to train using only information from smaller subgroups, we construct words in $S_{n}$ that permute at most $m$ elements and represent these as factorised expressions into general transpositions that are element of $S_{n}$. This is implemented by first generating a word in $S_{m}$ represented by a tuple $(\mathbf{i},\mathbf{j})=(i_{1},j_{1},\ldots,i_{n-1},j_{n-1})\in\{1,\ldots,m\}^{2(n-1)}$ according to its factorisation. [^1] For implementation purposes we generate here unreduced words of length $n-1$ which is larger than they need to be for $S_{m}$. We then convert $(\mathbf{i},\mathbf{j})$ to a tuple corresponding to a word in the larger group $S_{n}$ by relabeling using the map
@@ 3.2 Adjacent transpositions
This principle was implemented in a slightly different way for compatibility reasons. First, we generate a word in $S_{m}$ of length $N=n(n-1)/2$. Let $\mathbf{x}=(x_{1},\dots,x_{N})$ represent the input form of this word. Then, we generate a composition $\mu=(\mu_{1},\dots,\mu_{k})$ where $\mu_{1}+\dots+\mu_{k}=m$, representing the possible window sizes, and choose a corresponding set of admissible offsets $O=\{o_{1},\ldots,o_{k}\}$. Finally, we generate a set of integers $\{\ell_{1},\ldots,\ell_{N}\}$ uniformly at random with $\ell_{i}\in\{1,\ldots,k\}$, and perform the map
@@
InputTokenembedding$+$PositionalencodingMulti-head attentionAdd+normFeed forwardAdd+norm×$\times N_{L}$inearSoftmaxOutput probabilities
@@ 5 Results
Figure 3: Error and loss for training and validation using adjacent transpositions.
@@ 5.1 Performance
error=#incorrect predictions#test $\text{error}=\frac{\#\text{incorrect predictions}}{\#\text{test datapoints}}.$ (11)
For general transpositions we train on $S_{10}$ and test on $S_{25}$. We obtained zero training loss and near perfect test performance as shown in Figure 2. For adjacent transpositions, where we expect the complexity to be harder BayerDiaconis1992 ; Lacoin2016 ), we tested on $S_{16}$ and obtained similar performance, see Figure 3.
@@ 5.2 Interpretation
We have not observed grokking even though validation accuracy is optimal. Figures 4 and 5 depict heatmaps of the self-similarity matrices for the token embedding $M_{T}^{(E)}$ and position embedding $E_{P}$ defined in section 4.2. Let $A$ denote $E_{P}$ with normalized rows, then the matrix depicted in the image is given by $AA^{T}$.
@@
(b) Position embedding for $n=25$. The first $n-1$ rows and columns correspond to positions of transposition tokens and the last $n$ rows and colums to positions of permutation tokens.
@@
(a) Token embedding for $n=16$. The first $n-1$ rows and columns correspond to transposition tokens, the last $n$ rows and columns to permutation tokens.
@@
(b) Position embedding for $n=16$. The first $n(n-1)/2$ rows and columns correspond to positions of transposition tokens and the last $n$ rows and colums to positions of permutation tokens.

### v2

@@ 1.1 Related work on symmetric group
Under continued training we assume that the network generalizes, i.e. gradually shifts from memorization to learning an underlying algorithm. While we achieve test accuracy of nearly 100%, at the time of writing it is an open question whether our models have discovered a general algorithm or learned specific representations.
Interpretability of group multiplication in small permutation groups has been studied in a series of works ChughtaiCN2021 ; StanderYFB2023 ; Wu2024 (see also NandaCLSS2023 ). Analysis of symmetric group models using mechanistic interpretability can be done using different circuit level approaches Olah2020 ; Elhage2021 ; Olsson2022 ; Quirke2024 ; Zhang2023 ; Zhong2023 ; Gross2024 . Examining activations of individual neurons and intervening in various parts of the network can help us to discover circuits that implement the algorithm. We aim to perform a similar analysis of our models in a future work.
In contrast with previous works on learning tasks related to the symmetric group, we investigate whether a model is able to generalize from training on $S_{m}$ data to testing on $S_{n}$ data with $m<n$. This is out of distribution (OOD) learning which in the context of mathematics learning tasks has also been studied in Charton2022 . We test OOD learning in the context of the symmetric group, using a scalable presentation with general transpositions as group generators, as well as a less scalable, more local approach using only elementary transpositions. One motivation for this study is to identify an effective machine learning approach to improve on the unknot problem GukovHRS:2020 .
@@ 3 What are we learning?
The experiment reported in this note investigates whether a transformer can learn the symmetric group $S_{n}$ by training only on subgroups isomorphic to $S_{m}$ with $m<n$. That is, given a word $w\in S_{n}$ can the corresponding permutation $(\sigma(1),\ldots,\sigma(n)):=(1,\ldots,n)\cdot w$ be correctly predicted by a transformer without explicitly hardcoding the group relations (4) and training only on permutations that do not permute more than $m<n$ elements.
@@ 3.1 General transpositions
$\hskip-170.71652pt\Phi:\mathbf{x}\mapsto\mathbf{p},$
@@
In order to train using only information from smaller subgroups, we construct words in $S_{n}$ that permute at most $m$ elements and represent these as factorised expressions into general transpositions that are element of $S_{n}$. This is implemented by first generating a word in $S_{m}$ represented by a tuple $(\mathbf{i},\mathbf{j})=(i_{1},j_{1},\ldots,i_{n-1},j_{n-1})\in\{1,\ldots,m\}^{2(n-1)}$ according to its factorisation. [^1] For implementation purposes we generate here unreduced words of length n1n-1 which is larger than they need to be for $S_{m}$. We then convert $(\mathbf{i},\mathbf{j})$ to a tuple corresponding to a word in the larger group $S_{n}$ by relabeling using the map
@@ 3.2 Adjacent transpositions
This principle was implemented in a slightly different way for compatibility reasons. First, we generate a word in $S_{m}$ of length N=n(n1)/2N=n(n-1)/2. Let $\mathbf{x}=(x_{1},\dots,x_{N})$ represent the input form of this word. Then, we generate a composition $\mu=(\mu_{1},\dots,\mu_{k})$ where $\mu_{1}+\dots+\mu_{k}=m$, representing the possible window sizes, and choose a corresponding set of admissible offsets $O=\{o_{1},\ldots,o_{k}\}$. Finally, we generate a set of integers $\{\ell_{1},\ldots,\ell_{N}\}$ uniformly at random with $\ell_{i}\in\{1,\ldots,k\}$, and perform the map
@@
[…]
@@ 5 Results
Figure 3: Single token error and loss for training and validation using adjacent transpositions.
@@ 5.1 Performance
$\text{error}=\frac{\#\text{incorrect predictions}}{\#\text{test datapoints}}.$ (11)
For general transpositions we train on $S_{10}$ and test on $S_{25}$. We obtained zero training loss and near perfect test performance as shown in Figure 2. For adjacent transpositions, where we expect the complexity to be harder BayerDiaconis1992 ; Lacoin2016 ), we tested on $S_{16}$ and obtained similar performance, see Figure 3. We elaborate on validation and test data in Section A.4.
@@ 5.2 Interpretation
Figures 4 and 5 depict heatmaps of the self-similarity matrices for the token embedding $M_{T}^{(E)}$ and position embedding $E_{P}$ defined in section 4.2. Let $A$ denote $E_{P}$ with normalized rows, then the matrix depicted in the image is given by $AA^{T}$.
@@
(b) Position embedding for $n=25$. The first n1n-1 rows and columns correspond to positions of transposition tokens and the last $n$ rows and colums to positions of permutation tokens.
@@
(a) Token embedding for $n=16$. The first n1n-1 rows and columns correspond to transposition tokens, the last $n$ rows and columns to permutation tokens.
@@
(b) Position embedding for $n=16$. The first n(n1)/2n(n-1)/2 rows and columns correspond to positions of transposition tokens and the last $n$ rows and colums to positions of permutation tokens.
@@ Acknowledgments
Appendix A Appendix: Technical information
A.1 Model hyperparameters
All runs were trained using AdamW and cross entropy loss. The model is a standard transformer architecture with masked multi-headed attention. A custom attention mask was used, which is discussed later in the appendix. During training, we used a reduce-on-plateau learning rate scheduler with a reduction factor of 0.1 and a patience of 10 epochs. Teacher forcing was used during training.
We used the following hyperparameters for our runs.
Transpositions Dataset size Context length Vocabulary size
General 8,000,000 50 652
Elementary 16,000,000 136 34
Learning rate Weight decay Batch size Embedding dimension
0.0003 0.05 1024 402
0.0003 0.05 1024 402
Head count Block count FP format Trainable parameters
6 5 bf16 10,261,452
6 5 bf16 9,799,152
Here block count refers to $N_{L}$ as depicted in figure 1. All the code can be found on our GitHub Petschack_github  [^3] Transformer code is in the scaling-generator folder. Data generation code is in the fast-data-gen folder..
A.2 Custom attention mask
We would like all the tokens that make up the word to be able to attend to each other, but we also wanted to make sure that the model could not attend to future autoregression tokens. As such, we created the custom attention mask
$\left[\begin{array}[]{c|c}\mathbf{1}_{N\times N}&\mathbf{0}_{N\times n}\\ \hline\cr\mathbf{1}_{n\times N}&L_{n\times n}\end{array}\right],$ (12)
where $L$ denotes a lower triangular matrix.
A.3 Architecture details
This section elaborates on the architecture depicted in figure 1. The feed forward block has the following structure:
$\text{input}\to\text{Affine}(D,4D)\to\operatorname{ReLU}\to\text{Affine}(4D,D)\to\text{output}.$
Here $\text{Affine}(A,B)$ denotes an affine transformation from $A$ to $B$ dimensions, ie. a PyTorch linear layer. The final decoding layer (purple) is a $\text{Linear}(D,T)$ non-affine linear transformation. This is slightly inefficient as only a small subset of the tokens are valid outputs, but it does not make a large difference in practice.
A.4 Validation and test sets.
The validation set is size $10^{5}$ and is drawn from the same distribution as the training data, i.e. it is restricted to permutations of $m<n$ elements. The test data has size $10^{6}$ and is drawn from all possible permutations. The model for general transpositions found the correct permutation with 100% accuracy and the case of elementary transpositions achieved 99.993% accuracy. All the datasets are generated independently and there is no specific check against shared elements, but in practice the sample space is so large that it never happened (this was checked several times).
When generating random words, it occasionally happens that the final permutation lies within the smaller subgroup by chance. To make sure this did not effect our results, we calculated an upper bound on the amount of test data this applied to by determining which fraction of the test words permuted at most $m$ elements. For the case of general transpositions this was 0.7315%, and for the case of adjacent transpositions it was 2.243%, a relatively small percentage of the data in both cases.